# Accelerated Bayesian Inference for Pulsar Timing Arrays: Normalizing Flows for Rapid Model Comparison Across Stochastic Gravitational-Wave Background Sources

### v1

@@ Abstract
The recent detection of nanohertz stochastic gravitational-wave backgrounds (SGWBs) by pulsar timing arrays (PTAs) promises unique insights into astrophysical and cosmological origins. However, traditional Markov Chain Monte Carlo (MCMC) approaches become prohibitively expensive for large datasets. We employ a normalizing flow (NF)-based machine learning framework to accelerate Bayesian inference in PTA analyses. For the first time, we perform Bayesian model comparison across SGWB source models in the framework of machine learning by training NF architectures on the PTA dataset (NANOGrav 15-year) and enabling direct evidence estimation via learned harmonic mean estimators. Our examples include 10 conventional SGWB source models such as supermassive black hole binaries, power-law spectrum, cosmic strings, domain walls, scalar-induced GWs, first-order phase transitions, and dual scenario/inflationary gravitational wave. Our approach jointly infers 20 red noise parameters and 2 SGWB parameters per model in $\sim 20$ hours (including training), compared to $\sim 10$ days with MCMC. Critically, the NF method preserves rigorous model selection accuracy, with small Hellinger distances ($\lesssim 0.3$) relative to MCMC posteriors, and reproduces MCMC-based Bayes factors across all tested scenarios. This scalable technique for SGWB source comparison will be essential for future PTA expansions and next-generation arrays such as the SKA, offering orders-of-magnitude efficiency gains without sacrificing physical interpretability.
@@ I Introduction
To address this challenge, we build on recent developments in machine learning by implementing a normalizing flow (NF)-based Bayesian inference pipeline [14, 15]. Our architecture is trained on forward-simulated pulsar timing residuals for multiple SGWB+noise models, including realistic HD correlations and red noise components (see Appendix B for the workflow of our training). The NF model maps between the parameter space and a uniform latent distribution via invertible autoregressive flows, enabling efficient posterior reconstruction for each model.
Crucially, we show that this framework not only replicates MCMC-level accuracy for parameter inference, but also enables direct estimation of model evidence through a learned harmonic mean estimator (HME) [16, 17, 18]. Applied to 10-pulsar subsets of NG15 data, our pipeline yields robust posterior distributions and Bayes factors for ten SGWB source models, including variations of dual inflationary/bouncing universe scenarios. Compared to traditional inference workflows, our method reduces runtime by an order of magnitude (from $\sim$ 10 days to $\sim$ 20 hours, see Appendix F for comparative timing of NF and MCMC methods), while maintaining physical interpretability and consistency with MCMC benchmarks (Hellinger distances $\lesssim 0.3$).
@@ II Extracting the SGWB Power Spectrum from Pulsar Timing Residuals
with $\langle a_{I}(f)b_{J}(f^{\prime})\rangle=0$ and
@@
where the red noise $r_{I}^{\mathrm{RN}}(t)$ and SGWB $r_{I}^{\mathrm{SGWB}}(t)$ are captured by a single power spectral density (PSD) matrix,
@@ III.1 Normalizing Flow Model Construction and Training Process
Training NF architectures aims to optimize the probability density $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$ for simulated parameter vectors $\tilde{\theta}=\{\tilde{\theta}_{Di}^{(j)}\}$ under the physical model =$\mathcal{H}=\{\mathcal{H}^{(j)}\}$ (encompassing both noise and SGWB). Here, $\tilde{\mathbf{x}}=\{\tilde{\mathbf{x}}_{i}^{(j)}\}$ denotes the simulated timing residuals, with $i=1,\dots,2\times 10^{5}$ (the size of the training set), $j=1,\dots,10$ (the number of SGWB models in this study), $D=22$ (the dimensionality of each SGWB+noise parameter set, comprising 20 noise parameters ($2N_{\mathrm{pulsars}}$) plus 2 SGWB parameters), $\phi$ denotes the weight parameters of this NF-based machine learning model (the weight parameter file is saved after each training iteration, and posterior sampling is performed using this file along with the model script after convergence, $\phi\rightarrow\phi_{\mathrm{best}}$.). The workflow of the NF-based machine learning pipeline for SGWB analysis in this study is illustrated in Fig. 10 of Appendix B, which outlines data extraction from the NG15 dataset, generation of simulated dataset, NF model training, posterior inference of observational data, Bayes factors computation and SGWB model comparisions.
In NF-based ML, $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$ is mapped from a uniform base distribution $p_{\mathrm{base}}(\tilde{\textbf{z}}_{Di}^{(j)})=\text{Uniform}[-1,1]$ by the Jacobian determinant $\left|\det\left(\partial\tilde{\mathbf{z}}_{Di}^{(j)}/\partial\tilde{\theta}_{Di}^{(j)}\right)\right|$,
@@
where $\mathbf{\tilde{z}}\equiv T_{\phi}(\tilde{\theta};\tilde{\mathbf{x}},\mathcal{H})$, with $T_{\phi}$ an invertible mapping $T_{\phi}:\tilde{\theta}\mapsto\tilde{\mathbf{z}}$ built from autoregressive flows and permutation layers. This mapping transforms the simulated parameter vector $\tilde{\theta}$ into $\tilde{\mathbf{z}}$, which follows the base distribution $p_{\mathrm{base}}(\tilde{\textbf{z}})$. Here, $\phi$ denotes the machine-learning model’s weight parameters (saved to file after each training iteration). For each training iteration, the Jacobian determinant $\left|\det\left(\partial\tilde{\mathbf{z}}_{Di}^{(j)}/\partial\tilde{\theta}_{Di}^{(j)}\right)\right|$ tracks the change in probability density under the mapping.
@@
Before training of $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$, we firstly use the Python function np.random.uniform to generate simulated parameters $\tilde{\theta}=\{\tilde{\theta}_{Di}^{(j)}(\mathcal{H}^{(j)})\}$ for each SGWB+noise model $\mathcal{H}^{(j)}$ (sampled $2\times 10^{5}$ parameter points (22D: 20 red noise + 2 SGWB parameters) from the prior in Table 5) and Table 6) in Appendix D, and saves the sampling results to a file.
@@
Then we use get_rawdata.py to call micropta_SGWB$(j)$.py (this script is implemented based on the definitions in Eqs. 2 and varies for different SGWB models) to generate simulated timing residuals (4944-dimensional for the NG15 dataset) $\tilde{\mathbf{x}}=\{\tilde{\mathbf{x}}_{i}^{(j)}\!\left(\mathbf{d}_{\text{obs}},\tilde{\theta}_{Di}^{(j)},\mathcal{H}^{(j)}\right)\}$ from the parameter set $\tilde{\theta}=\{\tilde{\theta}_{Di}^{(j)}(\mathcal{H}^{(j)})\}$ with the pulsar observational data $\mathbf{d}_{\text{obs}}$ (including times of arrival and pulsar positions). Then we use get_rawdata.py to split the resulting dataset, $\tilde{\mathbf{x}}=\{\tilde{\mathbf{x}}_{i}^{(j)}\!\left(\mathbf{d}_{\text{obs}},\tilde{\theta}_{Di}^{(j)},\mathcal{H}^{(j)}\right)\}$, to be the training set and validation set (9:1). Note that in simulation, different SGWB+noise models $\mathcal{H}^{(j)}$ yield different simulated timing residuals $\tilde{\mathbf{x}}$, as the SGWB contribution depends on the specific model.
@@
To perform training, we use these two prepared simulated datasets, ($\left(\{\tilde{\theta}_{Di}^{(j)}\},\{\tilde{\mathbf{x}}_{i}^{(j)}\}\right)$ (stored in two separate files, simultaneously loaded during training), to opitimize the parameter probability density $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$ according to Eq. (8). In particular, during each training iteration, the model loads pairs of simulated parameters $\tilde{\theta}$ and residuals $\tilde{\mathbf{x}}$, then updates the weight parameters $\phi$ to minimize the loss function (the negative log-likelihood), thereby maximizing the model likelihood. Following Ref. [21] (Eq. (14)), the loss function is defined as:
@@
where $\alpha$ is the learning rate, and $n$ indexes the training iterations (with $n=0$ referring to the initial state of the weights, $\phi_{\mathrm{init}}$, which are randomly initialized according to PyTorch defaults and normal distributions; in this study, our result is convengent around $n=50$.). the validation set’s simulated residual files and corresponding parameter files are employed to recompute Eq. (III.1), yielding the training loss values that serve as a standard for evaluating model performance.
@@ III.2 Posterior Parameter Inference of Observational Data with Trained Normalizing Flows
The invertible mapping $T_{\phi_{\mathrm{best}}}:\theta\mapsto\mathbf{z}$ combines autoregressive flows and permutation layers [21], enabling rapid inverse transformations for efficient sampling, $T_{\phi_{\text{best}}}^{-1}$, in this study. More specifically, our approach employing NF-based ML method jointly infers 20 red noise parameters and 2 SGWB parameters per model in $\sim 20$ hours (including training), compared to $\sim 10$ days with MCMC. For more details of timing comparision between NF and MCMC method, see Appendix F.
@@ IV Visualization of posterior distribution from NF-based ML
In Fig. 1-Fig. 9, we illustrate the result from the NF and the reweighted NF, and compare them with the result from MCMC (benchmark method). It is evident that they agree well in the high-density regions, indicating that the NF and the reweighted NF methods effectively captures the main parameter constraints from MCMC method for various SGWB source models. In particular, both the 1-$\sigma$ and 2-$\sigma$ range from the NF analysis cover (is broader than) the corresponding regions from MCMC, suggesting that the NF method adopts a more conservative coverage in the tails. This behavior is likely related to the chosen training epochs ($n=50$) and number of analyzed pulsars ($N_{\mathrm{pulsars}}=10$ for NL while $N_{\mathrm{pulsars}}=68$ for MCMC) in this study. However, these differences do not affect the characterization of the core posterior structure; the central estimates from both methods essentially overlap, demonstrating that the NF approach can achieve comparable accuracy to traditional MCMC while significantly enhancing computational efficiency.
@@
Combing these reweighted posterior distributions with the Hellinger distance $H$ calculations in Table 1 demonstrates that, after reweighting, the posterior distribution more closely matches the MCMC-derived posterior than the direct NF sampling results. In particular, the Hellinger distance is bounded between 0 and 1, with smaller values indicating closer agreement between the distributions. In practice, $H<0.3$ implies that the two distributions are well aligned. For more details of Hellinger distance calculations see Appendix H.
@@ V Bayes Factor and SGWB source model comparisons
where $\varphi\left(\theta^{(j)}_{(D=2)_{l},\text{SGWB}}\right)$ is an arbitrary chosen normalized density introduced to remedy the exploding variance problem of original HME [28]. Specifically, we employ the Python package harmonic [18] with the two-dimensional SGWB parameter samples $\theta^{(j)}_{(D=2)_{l},\mathrm{SGWB}}$ and their corresponding likelihoods $\mathcal{L}\bigl{(}\mathbf{x}_{\mathrm{obs}}\mid\theta^{(j)}_{(D=2)_{l},\mathrm{SGWB}},\mathcal{H}^{(j)}\bigr{)}$ to compute the evidence $Z^{(j)}$ for each SGWB source model. The Bayes factors $\mathrm{BF}_{ij}=\frac{Z^{(i)}}{Z^{(j)}}$ are then listed in Table 3. For physical intepretation of Table 3, see Appendix I.
For the NF results in Table 3, each model was trained on a dataset of $2\times 10^{5}$ samples for 50 epochs using identical hyperparameters, and we selected the checkpoint with the lowest loss near epoch 50 for posterior sampling. To reduce the variance of the learned HME, we discarded the lowest 10% of likelihood values when computing both the log-likelihood and the log-evidence. In the same table, we also report Bayes factors derived from evidence estimates obtained via Nested Sampling [26] applied to the MCMC posterior samples.
@@
Table 3 presents a comprehensive comparison of Bayes factors across all SGWB source models (see Appendix C for model descriptions). In most cases, the NF‐derived Bayes factors agree with those from MCMC, with NF values lying within the uncertainties of traditional nested‐sampler estimates. Only a few models show minor discrepancies, likely due to variations in flow‐model training quality and finite training data. This concordance—together with the Hellinger distances reported in Table 1—demonstrates that rapid SGWB source model comparison can be achieved in an NF‐based ML framework without sacrificing accuracy. Our results pave the way for efficient SGWB source discrimination in future PTA expansions and next‐generation arrays such as the SKA, offering orders‐of‐magnitude gains in computational efficiency while preserving physical interpretability.

### v2

@@ Abstract
The recent detection of nanohertz stochastic gravitational-wave backgrounds (SGWBs) by pulsar timing arrays (PTAs) promises unique insights into astrophysical and cosmological origins. However, traditional Markov Chain Monte Carlo (MCMC) approaches become prohibitively expensive for large datasets. We employ a normalizing flow (NF)-based machine learning framework to accelerate Bayesian inference in PTA analyses. For the first time, we perform Bayesian model comparison across SGWB source models in the framework of machine learning by training NF architectures on the PTA dataset (NANOGrav 15-year) and enabling direct evidence estimation via learned harmonic mean estimators. Our examples include 10 conventional SGWB source models such as supermassive black hole binaries, power-law spectrum, cosmic strings, domain walls, scalar-induced GWs, first-order phase transitions, and dual scenario/inflationary gravitational wave. Our approach jointly infers 20 red noise parameters (10 pulsars) and 2 SGWB parameters per model in $\sim 20$ hours (including training), compared to $\sim 10$ days with MCMC (68 pulsars). Critically, the NF method preserves rigorous model selection accuracy, with small Hellinger distances ($\lesssim 0.3$) relative to MCMC posteriors, and reproduces MCMC-based Bayes factors across all tested scenarios. This scalable technique for SGWB source comparison will be essential for future PTA expansions and next-generation arrays such as the SKA, may offer substantial efficiency gains without sacrificing physical interpretability.
@@ I Introduction
To address this challenge, we build on recent developments in machine learning by implementing a normalizing flow (NF)-based Bayesian inference pipeline [13, 14, 15]. Our architecture is trained on forward-simulated pulsar timing residuals for multiple SGWB+noise models, including realistic HD correlations and red noise components (see Appendix B for the workflow of our training). The NF model maps between the parameter space and a uniform latent distribution via invertible autoregressive flows, enabling efficient posterior reconstruction for each model.
Crucially, we show that this framework not only replicates MCMC-level accuracy for parameter inference, but also enables direct estimation of model evidence through a learned harmonic mean estimator (HME) [16, 17, 18]. Applied to 10-pulsar subsets of NG15 data, our pipeline yields robust posterior distributions and Bayes factors for ten SGWB source models, including variations of dual inflationary/bouncing universe scenarios. Compared to traditional inference workflows MCMC (68 pulsars), our method reduces runtime from $\sim$ 10 days to $\sim$ 20 hours (10 pulsars) (see Sec. VI for more detailed discussion on the scalability trend, and see Appendix F for comparative timing of NF and MCMC methods), while maintaining physical interpretability and consistency with MCMC benchmarks (Hellinger distances $\lesssim 0.3$).
@@ II Extracting the SGWB Power Spectrum from Pulsar Timing Residuals
with $\langle a_{I}(f)b_{J}(f^{\prime})\rangle=0$,
@@
and both $a_{I}$ and $b_{I}$ are drawn from the Gaussian distributions, where the red noise $r_{I}^{\mathrm{RN}}(t)$ and SGWB $r_{I}^{\mathrm{SGWB}}(t)$ are captured by a single power spectral density (PSD) matrix,
@@ III.1 Normalizing Flow Model Construction and Training Process
Training NF architectures aims to optimize the probability density $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$ for simulated parameter vectors $\tilde{\theta}=\{\tilde{\theta}_{Di}^{(j)}\}$ under the physical model =$\mathcal{H}=\{\mathcal{H}^{(j)}\}$ (encompassing both noise and SGWB). Here, $\tilde{\mathbf{x}}=\{\tilde{\mathbf{x}}_{i}^{(j)}\}$ denotes the simulated timing residuals, with $i=1,\dots,2\times 10^{5}$ (the size of the training set), $j=1,\dots,10$ (the number of SGWB models in this study), $D=22$ (the dimensionality of each SGWB+noise parameter set, comprising 20 noise parameters ($2N_{\mathrm{pulsars}}$) plus 2 SGWB parameters), $\phi$ denotes the weight parameters of this NF-based machine learning model (the weight parameter file is saved after each training iteration, and posterior sampling is performed using this file along with the model script after convergence, $\phi\rightarrow\phi_{\mathrm{best}}$.). The workflow of the NF-based machine learning pipeline for SGWB analysis in this study is illustrated in Fig. 11 of Appendix B, which outlines data extraction from the NG15 dataset, generation of simulated dataset, NF model training, posterior inference of observational data, Bayes factors computation and SGWB model comparisons.
In NF-based ML, $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$ is mapped from a uniform base distribution $p_{\mathrm{base}}(\tilde{\textbf{z}}_{Di}^{(j)})=\text{Uniform}[-1,1]$ (as we assume a uniform prior on the SGWB model parameters  [^1] The relation between the base distribution $\tilde{\theta}$ and the prior distribution $\theta$ is given by [30]: $\tilde{\theta}=2\,\frac{\theta-\theta_{\mathrm{min}}}{\theta_{\mathrm{max}}-\theta_{\mathrm{min}}}-1,$ (8) where $\theta_{\mathrm{max}}$ and $\theta_{\mathrm{min}}$ denote the upper and lower bounds of $\theta$, respectively. For a uniform prior, one obtains a uniform base. For other prior distributions, such as a standard normal, one obtains a corresponding standard normal base. ) by the Jacobian determinant $\left|\det\left(\partial\tilde{\mathbf{z}}_{Di}^{(j)}/\partial\tilde{\theta}_{Di}^{(j)}\right)\right|$,
@@
where $\mathbf{\tilde{z}}\equiv T_{\phi}(\tilde{\theta};\tilde{\mathbf{x}},\mathcal{H})$, with $T_{\phi}$ an invertible mapping $T_{\phi}:\tilde{\theta}\mapsto\tilde{\mathbf{z}}$ built from autoregressive flows and permutation layers [22, 23, 24, 25, 26]. This mapping transforms the simulated parameter vector $\tilde{\theta}$ into $\tilde{\mathbf{z}}$, which follows the base distribution $p_{\mathrm{base}}(\tilde{\textbf{z}})$. Here, $\phi$ denotes the machine-learning model’s weight parameters (saved to file after each training iteration). For each training iteration, the Jacobian determinant $\left|\det\left(\partial\tilde{\mathbf{z}}_{Di}^{(j)}/\partial\tilde{\theta}_{Di}^{(j)}\right)\right|$ tracks the change in probability density under the mapping.
@@
Before training of $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$, we firstly use the Python function np.random.uniform to generate simulated parameters $\tilde{\theta}=\{\tilde{\theta}_{Di}^{(j)}(\mathcal{H}^{(j)})\}$ for each SGWB+noise model $\mathcal{H}^{(j)}$ (sampled $2\times 10^{5}$ parameter points (22D: 20 red noise parameters for 10 pulsars + 2 SGWB parameters) [^2] Each training epoch takes approximately $15\,\mathrm{minutes}$ for 22D. For 34D (20 red-noise parameters for 10 pulsars + 14 SGWB parameters for testing), it takes approximately $15\,\mathrm{minutes}$ and $30\,\mathrm{seconds}$ per epoch. from the prior in Table 6) and Table 7) in Appendix D, and saves the sampling results to a file.
@@
Then we use get_rawdata.py to call micropta_SGWB$(j)$.py (this script is implemented based on the definitions in Eqs. 2 and varies for different SGWB models) to generate simulated timing residuals (4944-dimensional for the NG15 dataset, see Table 5.) $\tilde{\mathbf{x}}=\{\tilde{\mathbf{x}}_{i}^{(j)}\!\left(\mathbf{d}_{\text{obs}},\tilde{\theta}_{Di}^{(j)},\mathcal{H}^{(j)}\right)\}$ from the parameter set $\tilde{\theta}=\{\tilde{\theta}_{Di}^{(j)}(\mathcal{H}^{(j)})\}$ with the pulsar observational data $\mathbf{d}_{\text{obs}}$ (including times of arrival and pulsar positions). The input size of $T_{\phi}$ equals the dimension of the residuals $\mathbf{x}$, which is 4944 for the selected NG15 dataset. Then we use get_rawdata.py to split the resulting dataset, $\tilde{\mathbf{x}}=\{\tilde{\mathbf{x}}_{i}^{(j)}\!\left(\mathbf{d}_{\text{obs}},\tilde{\theta}_{Di}^{(j)},\mathcal{H}^{(j)}\right)\}$, to be the training set and validation set (9:1) [28]. Note that in simulation, different SGWB+noise models $\mathcal{H}^{(j)}$ yield different simulated timing residuals $\tilde{\mathbf{x}}$, as the SGWB contribution depends on the specific model.
@@
To perform training, we use these two prepared simulated datasets, ($\left(\{\tilde{\theta}_{Di}^{(j)}\},\{\tilde{\mathbf{x}}_{i}^{(j)}\}\right)$ (stored in two separate files, simultaneously loaded during training), to optimize the parameter probability density $p_{\phi}(\tilde{\theta}_{Di}^{(j)}\mid\tilde{\mathbf{x}}_{i}^{(j)},\mathcal{H}^{(j)})$ according to Eq. (9). In particular, during each training iteration, the model loads pairs of simulated parameters $\tilde{\theta}$ and residuals $\tilde{\mathbf{x}}$, then updates the weight parameters $\phi$ to minimize the loss function (the negative log-likelihood), thereby maximizing the model likelihood. Following Ref. [29] (Eq. (14)), the loss function is defined as:
@@
where $\alpha$ is the learning rate ($\alpha=1\times 10^{-3}$ is taken as the initial value and, during training, it automatically decreases into a half when the loss stop decreasing), and $n$ indexes the training iterations (with $n=0$ referring to the initial state of the weights, $\phi_{\mathrm{init}}$, which are randomly initialized according to PyTorch defaults and normal distributions; in this study, our result is convergent around $n=50$.). the validation set’s simulated residual files and corresponding parameter files are employed to recompute Eq. (III.1), yielding the training loss values that serve as a standard for evaluating model performance.
@@ III.2 Posterior Parameter Inference of Observational Data with Trained Normalizing Flows
The invertible mapping $T_{\phi_{\mathrm{best}}}:\theta\mapsto\mathbf{z}$ combines autoregressive flows and permutation layers [29], enabling rapid inverse transformations for efficient sampling, $T_{\phi_{\text{best}}}^{-1}$, in this study. More specifically, our approach employing NF-based ML method jointly infers 20 red noise parameters and 2 SGWB parameters per model in $\sim 20$ hours (mainly due to the training process, whereas the inverse mapping $T^{-1}_{\phi_{\mathrm{best}}}$ takes only $40$ seconds in this case), compared to $\sim 10$ days with MCMC. For more details of timing comparison between NF and MCMC method, see Appendix F.
@@ IV Visualization of posterior distribution from NF-based ML
In Fig. 1-Fig. 9, we illustrate the result from the NF and the reweighted NF, and compare them with the result from MCMC (benchmark method). It is evident that they agree well in the high-density regions, indicating that the NF and the reweighted NF methods effectively captures the main parameter constraints from MCMC method for various SGWB source models. In particular, both the 1-$\sigma$ and 2-$\sigma$ range from the NF analysis cover (is broader than) the corresponding regions from MCMC, suggesting that the NF method adopts a more conservative coverage in the tails. This behavior is likely related to the chosen training epochs ($n=50$) and number of analyzed pulsars ($N_{\mathrm{pulsars}}=10$ for NL while $N_{\mathrm{pulsars}}=68$ for MCMC) in this study (for different epochs ($n=75$), see Fig. 12 of Appendix E). To test scalability, we have run epochs for $N_{\mathrm{pulsars}}=(8,9,10,11,12,13,14,15)$ with number of time residuals data $N_{\mathrm{time-residual}}=(3797,4220,4944,5604,6228,6585,7063,7463)$. These results suggest that the per-pulsar processing time - assuming each pulsar has a similar number of time residuals data- may remain approximately constant and that the overall training time appears to grow roughly linearly for $8\leq N_{\mathrm{pulsars}}\leq 15$. For more detailed discussion on the scalability trend, see Sec. VI. However, these differences do not affect the characterization of the core posterior structure; the central estimates from both methods essentially overlap, demonstrating that the NF approach can achieve comparable accuracy to traditional MCMC while significantly enhancing computational efficiency.
@@
Combing these reweighted posterior distributions with the Hellinger distance $H$ calculations in Table 1 demonstrates that, after reweighting, the posterior distribution more closely matches the MCMC-derived posterior than the direct NF sampling results. In particular, the Hellinger distance is bounded between 0 and 1, with smaller values indicating closer agreement between the distributions. In practice, $H<0.3$ implies that the two distributions are well aligned. Note that $H<0.3$ is only an empirical criterion for “well” aligned taken in this study. For more statistical interpretation and astrophysical applications, see Refs.[34, 35]. For more details of Hellinger distance calculations see Appendix H.
@@ V Bayes Factor and SGWB source model comparisons
where $\varphi\left(\theta^{(j)}_{(D=2)_{l},\text{SGWB}}\right)$ is an arbitrary chosen normalized density introduced to remedy the exploding variance problem of original HME [38]. Specifically, we employ the Python package harmonic [18] with the two-dimensional SGWB parameter samples $\theta^{(j)}_{(D=2)_{l},\mathrm{SGWB}}$ and their corresponding likelihoods $\mathcal{L}\bigl{(}\mathbf{x}_{\mathrm{obs}}\mid\theta^{(j)}_{(D=2)_{l},\mathrm{SGWB}},\mathcal{H}^{(j)}\bigr{)}$ to compute the evidence $Z^{(j)}$ for each SGWB source model. The Bayes factors $\mathrm{BF}_{ij}=\frac{Z^{(i)}}{Z^{(j)}}$ are then listed in Table 3. For physical interpretation of Table 3, see Appendix I.
For the NF results in Table 3, each model was trained on a dataset of $2\times 10^{5}$ samples for 50 epochs using identical hyperparameters, and we selected the checkpoint with the lowest loss near epoch 50 for posterior sampling. To reduce the variance of the learned HME, we discarded the lowest 10% of likelihood values when computing both the log-likelihood and the log-evidence. In the same table, we also report Bayes factors computed from evidence estimates obtained by applying the isocontour integration algorithm of Nested Sampling [36] to a kernel density estimator (KDE) [39] of the posterior sampled via MCMC, as implemented in Ceffyl [40].
@@
Table 3 presents a comprehensive comparison of Bayes factors across all SGWB source models (see Appendix C for model descriptions). In most cases, the NF‐derived Bayes factors agree with those from MCMC, with NF values lying within the uncertainties of traditional nested‐sampler estimates. Only a few models show minor discrepancies, likely due to variations in flow‐model training quality and finite training data. This concordance—together with the Hellinger distances reported in Table 1—demonstrates that rapid SGWB source model comparison can be achieved in an NF‐based ML framework without sacrificing accuracy. Our results pave the way for efficient SGWB source discrimination in future PTA expansions and next‐generation arrays such as the SKA, may offer substantial gains in computational efficiency while preserving physical interpretability.